# Isometric rigidity of the Wasserstein space over the plane with the maximum metric

### v1

@@ 4. Final remarks and open questions
We think that the ideas developed in this paper can be used to prove isometric rigidity of Wasserstein spaces that are built over certain normed spaces. Recent studies of the structure of Wasserstein isometries feature interesting examples of both rigid and non-rigid Wasserstein spaces. Kloeckner showed in [12] that the quadratic Wasserstein space over $\mathbb{R}^{n}$ admits non-trivial isometries. As a recent result of Che, Galaz-García, Kerin, and Santos-Rodríguez demonstrates [7], non-trivial isometries show up even if the underlying normed space $X$ can be written as $H\times Y$, where $H$ is a Hilbert space and $Y$ is a finite-dimensional normed space. On the other hand, non-quadratic Wasserstein spaces over Hilbert spaces [9, 10], and Wasserstein spaces over spheres, tori, and Heisenberg groups [4, 11] turned out to be rigid in the last years. Based on these recent developments, it would be an intricate question to study the problem of rigidity of quadratic and non-quadratic Wasserstein spaces over general normed spaces.

### v3

@@ 4. Final remarks and open questions
We think that the ideas developed in this paper can be used to prove isometric rigidity of Wasserstein spaces that are built over certain normed spaces. Recent studies of the structure of Wasserstein isometries feature interesting examples of both rigid and non-rigid Wasserstein spaces. Kloeckner showed in [13] that the quadratic Wasserstein space over $\mathbb{R}^{n}$ admits non-trivial isometries. As a recent result of Che, Galaz-García, Kerin, and Santos-Rodríguez demonstrates [8], non-trivial isometries show up even if the underlying normed space $X$ can be written as $H\times Y$, where $H$ is a Hilbert space and $Y$ is a finite-dimensional normed space. Furthermore, by an application of a recent result of Balogh, Stöher, Titkos and Virosztek (see Theorem 1.1. in [4]) it follows that the Wasserstein space $\mathcal{W}_{1}(Q,d_{1})$, (where $d_{1}$ is the $\ell_{1}$- metric on $Q$) is non-rigid as it contains mass-splitting isometries. This result is in sharp contrast to our main result, Theorem 1. On the other hand, non-quadratic Wasserstein spaces over Hilbert spaces [10, 11], and Wasserstein spaces over spheres, tori, and Heisenberg groups [5, 12] turned out to be rigid in the last years. Based on these recent developments, it would be an intricate question to study the problem of rigidity of quadratic and non-quadratic Wasserstein spaces over general normed spaces.
@@
Acknowledgments
The authors wish to express their gratitude to Eric Ströher for pointing out a gap in one of our proofs and to the anonymous referee for numerous comments improving both content and presentation of the manuscript.
Declaration
The authors declare that they have no competing interests.